OpenStudy (anonymous):

Mathematical Induction. [Number Theory] 1^2+3^2+5^2+...(2n-1)^2= n(2n-1)(2n+1)/3

OpenStudy (anonymous):

I have the basis

OpenStudy (anonymous):

I just need to work on the assume, and prove of the it.

OpenStudy (anonymous):

Now i just need to prove it.

OpenStudy (badhi):

asume this is true for n=m then, $$\sum\limits_{i=1}^m (2m-1)^2=\frac{m(2m+1)(2m-1)}{3}$$ add [2(m+1)-1]^2 to both sides, $$\begin{align*}\sum\limits_{i=1}^m (2m-1)^2+[2(m+1)-1]^2&=\frac{m(2m+1)(2m-1)}{3}+[2(m+1)-1]^2\\ \sum\limits_{i=1}^{m+1} (2m-1)^2&=\frac{4m^3-m+12m^2+12m+3}{3}\\ &=\frac{(m+1)(4m^2+8m+3)}{3}\\ &=\frac{(m+1)(2m+3)(2m+1)}{3}\\ &=\frac{(m+1)[2(m+1)+1][2(m+1)-1)}{3} \end{align*}$$ So, it is true for n=m+1

OpenStudy (badhi):

little typo in left side. it should be, $$\sum\limits_{i=1}^{m}(2i-1)^2$$ not $$\sum\limits_{i=1}^{m}(2m-1)^2$$